# Uncovering thermodynamic origin of counterflow and coflow instabilities in miscible binary superfluids

### v1

@@ II Hydrodynamics of homogeneous binary superfluids
$\Rightarrow$ $(-i\tilde{\omega}+M(\bm{k})\cdot\chi^{-1})\delta O=0\,,$
where $\tilde{\omega}=\omega-\bm{k}\cdot\bm{v^{n}}$. Then the spectrum of collective modes can be obtained by solving $\det(-i\tilde{\omega}+M\cdot\chi^{-1})=0$. The onset of a dynamical instability occurs when the imaginary part of $\tilde{\omega}$ becomes positive. This can happen when $\det(M\cdot\chi^{-1})=\det(M)/\det(\chi)$ vanishes and changes sign. It can be shown that $\det(M)>0$ following Goutéraux and Mefford (2024). Therefore, an instability can only occur when $\det(\chi)$ diverges and changes sign. Such instability from the divergence of the susceptibility matrix is obviously thermodynamic.
We now proceed by choosing a frame where $\bm{v^{n}}=0$. This is the case for holographic superfluids in the probe limit, and also for the superfluids modeled by GPE. We are interested in the effects from superfluid velocities, so we keep the temperature fixed from now on. For simplicity, we shall consider the case where the chemical potentials and charge densities are the same for both superfluid components. [^3] This is equivalent to setting the imbalance charge density $n_{-}=n_{1}-n_{2}$ identically to zero. This can straightforwardly be relaxed if needed. Then the thermodynamic quantities and sources we need to consider are just $\{n_{I}=(n,n),\,\bm{v^{s}_{I}}\}$ and $\{\mu_{I}=(\mu,\mu),\,\bm{\bar{h}_{I}}\}$. In this case, the EoMs are
@@
After considering the Onsager reciprocity and the positivity of entropy production, the first-order derivative corrections can be parameterized as
$\tilde{j^{i}}=$ $\sigma_{1}\partial^{i}\mu+\sigma_{2}\frac{v_{J}^{si}v_{J}^{sj}}{|v^{s}_{J}|^{2}}\partial_{j}\mu+\beta v_{I}^{si}\partial_{j}\bar{h}^{Ij},$ (10)
$\tilde{\mu_{I}}=$ $\beta v_{I}^{sj}\partial_{j}\mu+\delta_{1}\partial_{j}\bar{h}_{I}^{j}+\delta_{2}v_{I}^{si}v^{s}_{Ji}\partial_{j}\bar{h}^{Jj}\,,$
where $(\sigma_{1},\sigma_{2},\beta,\delta_{1},\delta_{2})$ are five non-hydrostatic positive parameters.  [^4] In principle, $\tilde{j^{i}}$ can be different for each component, i.e. $\Delta\tilde{j^{i}}=\tilde{j^{i}_{1}}-\tilde{j^{i}_{2}}\neq 0$. Nevertheless, from (1) one has $\partial_{i}\Delta\tilde{j^{i}}=0$ since $n_{1}=n_{2}$. On the other hand, only divergence of $\tilde{j^{i}}$ appears in the EoMs (9). Therefore, $\Delta\tilde{j^{i}}$ has no effects and one can simply take $\tilde{j^{i}_{1}}=\tilde{j^{i}_{2}}=\tilde{j^{i}}$.
The case with general superfluid velocities is still very complicated. We focus on the collinear case where $\bm{v^{s}_{1}}//\bm{v^{s}_{2}}//\bm{k}$ [^5] There is no conceptual obstacle generalising to other cases. But the resulting expressions are very lengthy. For later convenience, we change variables to
@@
By linearizing the EoMs (9) one can find the matrix $M_{AB}$ to be
@@
with $\sigma=\sigma_{1}+2\sigma_{2}$. In above equation, we have treated $\bar{h}_{I}$ as source for the convenience of linearizing EoMs. In practice, it is more convenient to treat $v^{s}_{I}$ as source. Let us denote $(n,v^{s}_{I})^{T}=\tilde{\chi}\cdot(\mu,\bar{h}_{I})^{T}$ and $(n,\bar{h}_{I})^{T}=\chi\cdot(\mu,v^{s}_{I})^{T}$. Then we can express $\tilde{\chi}$ in terms of $\chi$:
$\tilde{\chi}=\left(\begin{array}[]{ccc}\chi_{nn}+\chi_{n+}\frac{\chi_{n+}\chi_{--}-\chi_{n-}\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}+\chi_{n-}\frac{-\chi_{n+}\chi_{+-}+\chi_{n-}\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{\chi_{n+}\chi_{--}-\chi_{n-}\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{-\chi_{n+}\chi_{+-}+\chi_{n-}\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}\\ \frac{\chi_{n+}\chi_{--}-\chi_{n-}\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{\chi_{--}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{-\chi_{+-}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}\\ \frac{-\chi_{n+}\chi_{+-}+\chi_{n-}\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{-\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}\end{array}\right),$ (13)
where $\chi_{nn}=\partial n/\partial\mu|_{v^{s}_{I}}$, $\chi_{nI}=\partial n/\partial v^{s}_{I}|_{\mu}$ and $\chi_{IJ}=\partial\bar{h}_{I}/\partial v^{s}_{J}|_{\mu}$, We also have $\chi_{+-}=\chi_{-+}$ from Onsager relation (and also from $\chi_{AB}\equiv\delta^{2}f/\delta s_{A}\delta s_{B}$) where $f$ is the free energy density. By solving (8) with (12) and (13), we find two sound modes and one diffusive mode:
$\begin{split}\omega_{\pm}&=v_{\pm}k+i\Gamma_{\pm}k^{2}+\mathcal{O}(k^{3})\,,\\ \omega_{0}&=-i\Gamma_{0}k^{2}+\mathcal{O}(k^{3})\,,\end{split}$ (14)
@@
$\Gamma_{0}$ $=\frac{(\chi_{++}\chi_{--}-\chi_{+-}^{2})(\delta_{2}(v^{s}_{-})^{2}+\delta_{1})}{4\chi_{++}}\,,$
@@
$+4\beta(v_{+}^{s}+v_{-}^{s})v_{\pm}(v_{+}^{s}(v_{\pm}\chi_{n+}-\chi_{++})+v_{-}^{s}(v_{\pm}\chi_{n-}-\chi_{+-})))\,,$
with $\delta_{1}$, $\delta_{2}$, $\beta$ and $\sigma$ positive constants.
The criterion for the onset of instability is given by $\mathrm{det}(\tilde{\chi})=-\chi_{nn}/(\chi_{++}\chi_{--}-\chi_{+-}^{2})$ diverging and changing sign. It gives
@@
For the counterflow case where $v^{s}_{+}=0$ and the coflow case where $v^{s}_{-}=0$, we have $\chi_{+-}=0$, since $\chi_{+-}=\delta^{2}f((v^{s}_{+})^{2},(v^{s}_{-})^{2})/(\delta v^{s}_{+}\delta v^{s}_{-})\sim v^{s}_{+}v^{s}_{-}$. Therefore, in both cases, the criterion for the onset of instability is given by $\chi_{++}=0$ or $\chi_{--}=0$. We will see later this is indeed the correct criteria for onset of counterflow and coflow instabilities in the holographic binary superfluid model. It also explains the counterflow instability found from GPE.
III Counterflow instability from GPE
@@
The development of quantum turbulence from two counter-propagating superfluids of miscible BECs has been studied using the mean-field GPE Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b). In the mean field approximation, the condensate wave functions are described by $\Psi_{J}(x,y,t)=\sqrt{n^{s}_{J}(x,y,t)}e^{i\theta_{J}}(\mathbf{r},t)$ where $n^{s}_{J}$ and $\theta_{J}$ are the particle density and phase of the $J$-th component of the binary superfluids. The latter is nothing but the Goldstone field $\varphi_{J}$ for the $J$-th superfluid component. The coupled GPEs are given as (capital Latin indices are not summed here)
$i\partial_{t}\Psi_{I}=(-\frac{1}{2m_{I}}\nabla^{2}-\mu_{I}+g_{I}|\Psi_{I}|^{2}+g_{IJ}|\Psi_{J}|^{2})\Psi_{I},$ (18)
$(I,J=1,2,\quad I\neq J)\,,$
which describes binary superfluids without dissipation and normal fluid component. Here $m_{I}$ is the mass of the $I$-th component, and $g_{1},g_{2}$ and $g_{12}$ are the coupling constants. To obtain stable miscible condensates, we consider the conditions $g_{12}<\sqrt{g_{1}g_{2}}$ and $g_{I}>0$.
It was shown that the countersuperflow becomes unstable and quantized vortices are nucleated as the relative velocity exceeds a critical value, leading to isotropic quantum turbulence consisting of two superflows. In contrast to the Landau instability steaming from thermodynamic instability, it was argued that the counterflow instability is purely an “internal” instability of the isolated system of two superfluids without any influence from the external environment. As a dynamic instability, the physical origin of the counterflow instability observed in the coupled GPEs is still unclear.
Interestingly, we show that the counterflow instability of the two-component BECs via (18) is indeed a thermodynamic instability in the sense det($\tilde{\chi}$) diverges. For the case where there are two $U(1)$ conserved charges, one can find this also happens when (17) is satisfied. To be more specific, it is given by
$\chi_{++}=n^{s}_{+}+v^{s}_{-}\partial_{v^{s}_{+}}n^{s}_{-}+N=0\,.$ (19)
@@
For simplicity, let’s consider the case where $m_{1}=m_{2}=m$, $g_{1}=g_{2}=g$ and $\mu_{1}=\mu_{2}=\mu$. From (18), the superfluid densities for stationary and uniform state can be found to be
@@
This is exactly the critical velocity derived from Bogoliubov–de Gennes analysis Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b), with $n^{s}_{+}=2n$. Therefore, the dynamical counterflow instability found by numerically solving the coupled GPEs is essentially thermodynamic in origin.

### v2

@@ II.2 Linear analysis and the instabilities
$\Rightarrow$ $(-i\tilde{\omega}+M(\bm{k})\cdot\bar{\chi}^{-1})\delta O=0\,$,
where $\tilde{\omega}=\omega-\bm{k}\cdot\bm{v^{n}}$. The fact that the matrix $M$ only depends on $\bf k$} follows from our choice of frame. We show this explicitly in equation (17) below.
Then the spectrum of collective modes can be obtained by solving $\det(-i\tilde{\omega}+M\cdot\bar{\chi}^{-1})=0$. The onset of a dynamical instability occurs when the imaginary part of $\tilde{\omega}$ becomes positive. In Goutéraux and Mefford (2024), a general proof is given that provided positivity of entropy is imposed $\Delta\geq 0$, this can only occur when an eigenvalue of $\bar{\chi}$, and so $\det(\bar{\chi})$ also, diverges and changes sign, revealing the thermodynamic origin of these instabilities. In this work, we restrict to their study in the ‘probe’ limit where normal velocity and temperature fluctuations are set to zero. Thus, only the charge conservations equations and Josephson equations need to be solved. We further simplify the analysis by restricting to choices of the superfluid velocities relevant for the coflow and counterflow instabilities, meaning that we take both velocities in the same direction with equal (coflow) or opposite (counterflow) orientations. Then, also taking the wavenumber $\bf k$} in the same direction as the superfluid velocities, the modes can be solved for analytically and the onset of the instabilities can be studied by direct inspection, without needing to resort to the general proof in Goutéraux and Mefford (2024). [^4] The proof in Goutéraux and Mefford (2024) is presented for single-component superfluids, and gives rigorous justification for earlier results obtained in Goutéraux et al. (2023); Areán et al. (2024). Its extension to binary superfluids is straightforward.
We now consider a frame where the normal fluid is at rest, $\bm{v^{n}}=0$. This is the case for holographic superfluids in the probe limit, and also for the superfluids modeled by GPE. For simplicity, we shall consider the case where the chemical potentials and charge densities are the same for both superfluid components. [^5] This is equivalent to setting the imbalance charge density $n_{-}=n_{1}-n_{2}$ identically to zero. This can straightforwardly be relaxed if needed. Then the thermodynamic quantities and sources we need to consider are just $\{n_{I}=(n,n),\,\bm{v^{s}_{I}}\}$ and $\{\mu_{I}=(\mu,\mu),\,\bm{\bar{h}_{I}}\}$. In this case, the EoMs are
@@
Positivity of entropy production $\Delta\geq 0$ leads us to consider non-ideal constitutive relations in the form
$\left(\begin{array}[]{c}\tilde{j}^{i}\\ \tilde{\mu}_{I}\end{array}\right)=-\tilde{\Sigma}\cdot\left(\begin{array}[]{c}\partial_{j}\mu\\ \partial_{j}\bar{h}_{J}^{j}\end{array}\right)$ (11)
so that
$\Delta=\left(\partial_{i}\mu,\,\partial_{j}\bar{h}_{I}^{j}\right)\cdot\tilde{\Sigma}\cdot\left(\begin{array}[]{c}\partial_{j}\mu\\ \partial_{j}\bar{h}_{J}^{j}\end{array}\right)\geq 0$ (12)
means imposing that the matrix $\tilde{\Sigma}$ is positive-definite. It suffices then to parametrize $\tilde{\Sigma}$ in the most general way compatible with the index structures and with time reversal symmetry (Onsager reciprocity). [^6] Under time-reversal, $\left\{n,\mu,{\bf v_{s}^{I}},{\bf\bar{h}^{I}}\right\}\mapsto\left\{n,\mu,-\bf v_{s}^{I},-{\bf\bar{h}^{I}}\right\}$. Onsager symmetry then implies $G^{R}_{AB}(\omega,{\bf k},{\bf v_{s}^{I}})=\eta_{A}\eta_{B}G^{R}_{AB}(\omega,-{\bf k},-{\bf v_{s}^{I}})$, with $\eta_{A}$ the eigenvalue under time reversal of operator $A$. These constraints must also be obeyed by the static susceptibility matrix $\bar{\chi}_{AB}$ as well as by the matrix $\tilde{\Sigma}$. This implies in particular that the same coefficient $\beta$ appears in the off-diagonal block of this matrix. This gives:
$\tilde{\Sigma}=\left(\begin{array}[]{cc}\sigma_{1}\delta^{ij}+\sigma_{2}\frac{v_{J}^{si}v_{J}^{sj}}{|v_{J}^{s}|^{2}}&\beta v_{J}^{si}\\ \beta v_{I}^{sj}&\delta_{1}\delta_{IJ}+\delta_{2}v_{I}^{si}v_{Ji}^{s}\end{array}\right)$ (13)
where $(\sigma_{1},\sigma_{2},\beta,\delta_{1},\delta_{2})$ are five independent transport coefficients. [^7] In principle, $\tilde{j^{i}}$ can be different for each component, i.e. $\Delta\tilde{j^{i}}=\tilde{j^{i}_{1}}-\tilde{j^{i}_{2}}\neq 0$. Nevertheless, from (2) one has $\partial_{i}\Delta\tilde{j^{i}}=0$ since $n_{1}=n_{2}$. On the other hand, only divergence of $\tilde{j^{i}}$ appears in the EoMs (10). Therefore, $\Delta\tilde{j^{i}}$ has no effects and one can simply take $\tilde{j^{i}_{1}}=\tilde{j^{i}_{2}}=\tilde{j^{i}}$. We can now write $\Delta$ as:
$\Delta=$ $\,\sigma_{1}\left(\partial^{i}\mu+\frac{\beta}{4\sigma_{1}}v^{si}_{I}\partial^{k}\bar{h}_{kI}\right)^{2}$
$+\left(\delta_{2}-\frac{\beta^{2}}{2\sigma_{1}}\right)v^{si}_{I}\partial^{k}\bar{h}_{kI}v^{s}_{iJ}\partial^{j}\bar{h}_{jJ}$
$+\sigma_{2}\left(v^{si}_{I}\partial_{i}\mu+\frac{\beta|v^{s}_{J}|^{2}}{2\sigma_{2}}\partial_{k}\bar{h}^{k}_{I}\right)^{2}$
$+\left(\delta_{1}-\frac{\beta^{2}|v^{s}_{J}|^{2}}{4\sigma_{2}}\right)\partial_{k}\bar{h}^{k}_{I}\partial^{j}\bar{h}^{I}_{j}\,$, (14)
and imposing that it is a positive-definite quadratic form for any configuration of sources $\mu$, $\bar{h}_{I}^{i}$ results in the following constraints:
$\sigma_{1,2}\geq 0\,,\quad\beta^{2}\leq 4\sigma_{1}\delta_{2}\,,\quad\beta^{2}\leq\frac{4\sigma_{2}\delta_{1}}{|v^{s}_{J}|^{2}}\,.$ (15)
The case with general superfluid velocities is still very complicated. We focus on the collinear case where $\bm{v^{s}_{1}}//\bm{v^{s}_{2}}//\bm{k}$. [^8] There is no conceptual obstacle generalising to other cases. But the resulting expressions are very lengthy. For a single species superfluid, this is the sector where the dominant instability is found Amado et al. (2014); Schmitt (2015). Evidence that this continues to be true for binary superfluids was presented in Ishino et al. (2011). For later convenience, we change variables to
@@
By linearizing the EoMs (10), inserting the constitutive relations (11) and (13) we find the matrix $M_{AB}$ to be
@@
with $\sigma=\sigma_{1}+2\sigma_{2}$ (we present the linearized equations in the appendix F). In the equation above, we have treated $\bar{h}_{I}$ as a source as it makes linearizing the EoMs easier. In practice, it is more convenient to treat $v^{s}_{I}$ as a source (as this is easier to control the value of this parameter). Let us denote $(\delta n,\delta v^{s}_{I})^{T}=\bar{\chi}\cdot(\delta\mu,\delta\bar{h}_{I})^{T}$ and $(\delta n,\delta\bar{h}_{I})^{T}=\chi\cdot(\delta\mu,\delta v^{s}_{I})^{T}$. Then we can express $\bar{\chi}$ in terms of $\chi$:
$\bar{\chi}=\left(\begin{array}[]{ccc}\chi_{nn}+\chi_{n+}\frac{\chi_{n+}\chi_{--}-\chi_{n-}\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}+\chi_{n-}\frac{-\chi_{n+}\chi_{+-}+\chi_{n-}\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{\chi_{n+}\chi_{--}-\chi_{n-}\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{-\chi_{n+}\chi_{+-}+\chi_{n-}\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}\\ \frac{\chi_{n+}\chi_{--}-\chi_{n-}\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{\chi_{--}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{-\chi_{+-}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}\\ \frac{-\chi_{n+}\chi_{+-}+\chi_{n-}\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{-\chi_{-+}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}&\frac{\chi_{++}}{\chi_{++}\chi_{--}-\chi_{+-}\chi_{-+}}\end{array}\right),$ (18)
where $\chi_{nn}=\partial n/\partial\mu|_{v^{s}_{I}}$, $\chi_{nI}=\partial n/\partial v^{s}_{I}|_{\mu}$ and $\chi_{IJ}=\partial\bar{h}_{I}/\partial v^{s}_{J}|_{\mu}$, We also have $\chi_{+-}=\chi_{-+}$, Onsager reciprocity and as is clear from $\bar{\chi}_{AB}\equiv-\delta^{2}\bar{F}/\delta s_{A}\delta s_{B}$. By solving (9) with (17) and (18), we find two sound modes and one diffusive mode: [^9] In the more general case where the two U(1) densities are kept different, there are instead two pairs of superfluid sound modes Andreev and Bashkin (1975).
[…]
@@
$\Gamma_{0}$ $=-\frac{(\chi_{++}\chi_{--}-\chi_{+-}^{2})(\delta_{2}(v^{s}_{-})^{2}+\delta_{1})}{4\chi_{++}}\,$,
@@
$+4\beta(v_{+}^{s}+v_{-}^{s})v_{\pm}(v_{+}^{s}(v_{\pm}\chi_{n+}-\chi_{++})+v_{-}^{s}(v_{\pm}\chi_{n-}-\chi_{+-})))\,.$
Checking that $\Gamma_{\pm}<0$ provided the constraints (15) hold and that $\bar{\chi}$ is positive-definite is tedious, highlighting the interest of a more general derivation, Goutéraux and Mefford (2024).
The criterion for the onset of instability Goutéraux and Mefford (2024) is given by $\mathrm{det}(\bar{\chi})=-\chi_{nn}/(\chi_{++}\chi_{--}-\chi_{+-}^{2})$ diverging and changing sign. It gives [^10] In principle, it may also be that $\chi_{nn}$ changes sign. But this is not what we observe in the two microscropic examples based on the GPE and gauge/gravity duality.
@@
In the counterflow and coflow cases, the expressions for the modes become simpler and the analysis can be performed by direct inspection more easily.
In the counterflow case, $v^{s}_{+}=0$, we have $\chi_{+-}=0$, since $\chi_{+-}=\delta^{2}F\left[(v^{s}_{+})^{2},(v^{s}_{-})^{2}\right]/(\delta v^{s}_{+}\delta v^{s}_{-})\sim v^{s}_{+}v^{s}_{-}$. For a similar reason, we also have $\chi_{n+}=0$. Then, the expressions for the modes become
$\omega_{\pm}=$ $\pm\sqrt{\frac{\chi_{++}}{\chi_{nn}}}k-\frac{ik^{2}}{2\chi_{nn}}\left[\sigma-v_{-}\beta\chi_{n-}\right$.
$\left.+\frac{\chi_{n-}^{2}}{4}(\delta_{1}+v_{-}^{2}\delta_{2})+\frac{\chi_{nn}\chi_{++}}{4}\delta_{1}\right],$ (23)
$\omega_{0}=$ $-\frac{i}{4}k^{2}(\delta_{1}+v_{-}^{2}\delta_{2})\chi_{--}\,.$ (24)
It is clear from both expressions that an instability occurs if $\chi_{++}$ becomes negative, since then the sound speed becomes purely imaginary, with $\omega_{-}$ moving into the upper half plane. This is precisely what we observe in the next sections in the GPE Takeuchi et al. (2010b) and in holographic superfluids. In principle, $\chi_{--}$ changing sign would also cause an instability by sending $\omega_{0}$ into the upper half plane, but this is not what we find happens in the GPE and in holographic superfluids.
For the coflow case $v_{-}=0$, we have $\chi_{+-}=\chi_{n-}=0$. The modes simplify to
$\omega_{0}=$ $-i\frac{\chi_{--}}{4}\delta_{1}k^{2}\,$, (25)
$\omega_{\pm}=$ $v_{\pm}k+i\Gamma_{\pm}k^{2}\,$, (26)
$\Gamma_{\pm}=$ $-\frac{1}{2}v^{s}_{+}v_{\pm}\beta-\frac{(v_{+}^{s})^{2}}{8}(\chi_{++}-\chi_{n+}v_{\pm})\delta_{2}$
$\mp v_{\pm}\frac{16\sigma+\chi_{nn}^{2}(v_{+}-v_{-})^{2}\delta_{1}}{16\chi_{nn}(v_{+}-v_{-})}\,.$ (27)
The expression for $v_{\pm}$ is given in (20).
We also see two possible instabilities in this case, when either $\chi_{--}$ or $\chi_{++}$ change sign. In holographic superfluids, the former is realized first as $v_{+}^{s}$ is increased and leads to $\omega_{0}$ crossing to the upper half plane. A secondary instability occurs when $\chi_{++}$ changes sign, upon which both $v_{+}$ and $\Gamma_{+}$ vanish and then change sign, moving $\omega_{+}$ to the upper half plane. To highlight this, we expand the relevant dispersion relation around $\chi_{++}=0$, assuming $\chi_{n+}>0$: [^11] If $\chi_{n+}<0$, then $\omega_{-}$ becomes unstable instead.
$\omega_{+}=$ $\frac{\delta\chi_{++}}{2\chi_{n+}}k-ik^{2}\frac{\delta\chi_{++}}{16}\times$ (28)
$\left(\delta_{1}+\delta_{2}(v_{+}^{s})^{2}+\frac{4\beta v_{+}^{s}}{\chi_{n+}}+\frac{4\sigma}{\chi_{n+}^{2}}\right).$
When $\delta\chi_{++}<0$, the mode crosses to the upper half plane and becomes unstable.
III Counterflow instability from the GPE
@@
The development of quantum turbulence from two counter-propagating superfluids of miscible BECs has been studied experimentally and by using the mean-field GPE Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b); Beattie et al. (2013b); Yakimenko et al. (2013); Abad et al. (2014). In the mean field approximation, the condensate wave functions are described by $\Psi_{J}(\mathbf{r},t)=\sqrt{n^{s}_{J}(\mathbf{r},t)}e^{i\theta_{J}}(\mathbf{r},t)$ where $n^{s}_{J}$ and $\theta_{J}$ are the particle density and phase of the $J$-th component of the binary superfluids. The latter is nothing but the Goldstone field $\varphi_{J}$ for the $J$-th superfluid component. The coupled GPEs are given as (here, capital Latin indices are not summed over)
$i\partial_{t}\Psi_{I}=$ $\left(-\frac{1}{2m_{I}}\nabla^{2}-\mu_{I}+g_{I}|\Psi_{I}|^{2}+g_{IJ}|\Psi_{J}|^{2}\right)\Psi_{I},$ (29)
$\quad(I,J=1,2,\quad I\neq J)\,.$
They describe binary superfluids without dissipation or a normal fluid component. Here $m_{I}$ is the mass of the $I$-th component, and $g_{1},g_{2}$ and $g_{12}$ are the coupling constants. For more about BEC and GPE, readers can refer to the textbooks Pethick and Smith (2008); Pitaevskii and Stringari (2003). To obtain stable miscible condensates, we consider the conditions $g_{12}<\sqrt{g_{1}g_{2}}$ and $g_{I}>0$. It was shown that the countersuperflow becomes unstable and quantized vortices are nucleated as the relative velocity exceeds a critical value, leading to isotropic quantum turbulence consisting of two superflows.
We show that the counterflow instability of the two-component BECs via (29) continues to be given by the divergence of $\det(\bar{\chi})$. Since the derivatives involved in the criterion we derived assuming the existence of a hydrodynamic regime involve quantities that are well-defined even at zero temperature (superfluid velocity, charge densities), it is meaningful to evaluate them at zero temperature with the GPE. Further, perfect superfluid hydrodynamics can be derived from the GPE Pitaevskii and Stringari (2016); Sonin (2016), at zero temperature, neglecting dissipation and in the absence of a normal component. For the case where there are two $U(1)$ conserved charges, one can find $\det(\bar{\chi})$ diverges when (22) is satisfied. To be more specific, it is given by
$\chi_{++}=n^{s}_{+}+v^{s}_{-}\partial_{v^{s}_{+}}n^{s}_{-}+N=0\,$, (30)
@@
with $n^{s}_{+}=n^{s}_{1}+n^{s}_{2}$ and $n^{s}_{-}=n^{s}_{1}-n^{s}_{2}$. For simplicity, let’s consider the case where $m_{1}=m_{2}=m$, $g_{1}=g_{2}=g$ and $\mu_{1}=\mu_{2}=\mu$. From (29), the superfluid densities for stationary and uniform state can be found to be
@@
This is exactly the critical velocity derived from Bogoliubov–de Gennes analysis Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b), with $n^{s}_{+}=2n$. Therefore, we see that the microscopic Bogoliubov-de Gennes theory offers a microscopic realization of the macroscopic description of the dynamical counterflow instability that we established in (22). As we shall see now, the macroscopic criterion (22) applies to a completely different class of superfluids, which are those that have a gravitational dual.